OpenStudy (anonymous):

SIMPLIFY: X^2+9X-70 ---------- X-5

OpenStudy (anonymous):

answer is \[x+14\] factor and cancel

OpenStudy (anonymous):

THANKS VERY MUCH FOR THE HELP

OpenStudy (anonymous):

\[\frac{x^2+9x-70}{x-5}=\frac{(x+14)(x-5)}{x-5}=x+14\]

OpenStudy (anonymous):

yw